# MapReduce LoRA: Advancing the Pareto Front in Multi-Preference Optimization for Generative Models

### v4

@@ 1 Introduction
To address these limitations, we propose two complementary approaches—MapReduce LoRA and Reward-aware Token Embedding (RaTE)—that jointly advance multi-preference alignment. MapReduce LoRA decomposes multi-objective optimization into i) a Map phase that trains reward-specific LoRA experts in parallel, and ii) a Reduce phase that merges experts with user-controlled interpolation, folds the merged adapter into the base, and iterates to advance the Pareto front. Complementarily, RaTE introduces a lightweight inference-time control by assigning each reward a learned token embedding, enabling composable conditioning by appending multiple tokens to the input prompt. Together, MapReduce LoRA and RaTE enable efficient a posteriori customization without retraining, yielding unified generative models that excel across multiple reward dimensions.
Extensive experiments demonstrate that our proposed methods substantially advance the Pareto front of multi-preference alignment across both Text-to-Image and Text-to-Video generation (Fig. 1). On Stable Diffusion 3.5 Medium [3] and FLUX.1-dev [17], our method improves GenEval [11], PickScore [15], and OCR [6] by 36.1%, 4.6%, 55.7% and 32.7%, 4.3%, 67.1%, respectively. Beyond in-domain metrics, untargeted rewards—i.e., VQAScore [21], MPS [43], and VILA [14]—also improve by 1.85%, 6.49%, and 19.96%, validating the robustness and scalability of our method. On HunyuanVideo [16], visual and motion quality improve by 48.1% and 90.0%, achieving state-of-the-art results among post-trained diffusion systems. On language task, Helpful Assistant, with Llama-2 7B [38], helpful and harmless improve by 43.4% and 136.7%, respectively. Collectively, these results highlight the effectiveness of combining MapReduce LoRA and RaTE for efficient, controllable, and human-aligned generation. Our contributions are fourfold:
@@ 2.1 Flow-based Generative Models
Flow Matching (FM) [22] trains continuous normalizing flows by regressing the conditional velocity field along a path between the data distribution and a standard normal, yielding a score-free, simulation-free objective with more stable and efficient optimization than diffusion losses. Rectified Flow (RF) [25] specializes FM to a straight-line, approximately constant-speed path, rectifying trajectories and simplifying supervision, which improves stability and sampling efficiency. Recent Text-to-Image [8, 3, 17, 7] and Text-to-Video [16, 31, 39] methods adopt FM/RF to learn the velocity field directly, improving efficiency over diffusion-style denoising and often yielding stronger gradients for conditional generation.
@@ 2.2 Reinforcement Learning from Human Feedback (RLHF)
However, single-reward RLHF optimizes one preference dimension in isolation, while generative quality spans multiple and often conflicting objectives. This creates a natural bridge to Multi-Objective RL (MORL), where the objective is to balance conflicting rewards and approximate a Pareto-optimal solution—precisely the challenge in post-training generative models. MORL methods can be categorized into two classes: a priori [18, 1], which scalarizes objectives into a single signal (e.g., weighted sums), and a posteriori [33, 20], which approximate a Pareto set and enable test-time control. CaPO [18] calibrates and balances multiple T2I preferences, and MOPO [1] frames alignment as constrained, KL-regularized optimization; both are a priori and lack test-time control. Rewarded Soups [33] extends Linear Mode Connectivity [9, 28] to RL, showing linear weight soups can approximate Pareto fronts and reduce reward misspecification, yet multi-reward methods often underperform specialized single-reward experts. We therefore propose two a posteriori solutions, RaTE and MapReduce LoRA, to address reward conflict and achieve comparable performances to single-reward experts’ performances.
@@ 3.1 Preliminaries
where $A\in\mathbb{R}^{r\times k}$, $B\in\mathbb{R}^{d\times r}$, and $r\ll\min(d,k)$. During fine-tuning, $W$ remains fixed while only AA and $B$ are trained, optionally scaled by a factor $\alpha/r$; at inference, $\Delta W$ can be merged into $W$. Recent work [34] finds that, in common post-training settings (small/medium-scale SFT, reasoning, or RL), LoRA matches full fine-tuning performance when rank and layer coverage suffice. LoRA has been used for style [44] and skill control [20], but multi-preference control remains underexplored; we therefore study LoRA merging for multi-preference optimization.
@@
Pareto Fronts (PF) [30] represents the collection of non-dominated solutions, model parameters for which no other candidate can improve one objective without degrading another. Formally, $\mathrm{PF}=\left\{\theta\mid\nexists\,\theta^{\prime}\in\Theta,\;\{R_{i}(\theta^{\prime})\}_{i}\succ_{\mathbb{R}^{N}}\{R_{i}(\theta)\}_{i}\right\},$ where $\succ_{\mathbb{R}^{N}}$ denotes the dominance relation in the N-dimensional reward space indexed by $i$, and NN denotes the total number of reward objectives.
@@ 4.3 Quantitative Results
Reward-aware token embedding (RaTE). Tables 1 and 3 report RaTE performance on GenEval [11], PickScore [15], and OCR [6] using SD 3.5 M [3]. At inference, RaTE appends a trained control token to the end of the prompt based on the specified preference. RaTE enables per-reward control and jointly improves multiple rewards. In Table 1, token gains are orthogonal to transformer tuning: RaTE adds 4.20%, 1.03%, and 2.17% on GenEval [11], PickScore [15], and OCR [6], respectively. In Table 3, when using all three tokens, the first appended token accounts for most of the gain. Token control is effective on Stable Diffusion [3, 10], which uses explicit cross-attention for text conditioning. In contrast, FLUX.1-dev [17] performs joint text–image sequence modeling, causing a modified token embedding to perturb both text and image tokens across layers, which makes reward-specific token control substantially less stable. We leave this for future exploration.
@@
4.4 Ablation Study
[MapReduce LoRA] Effect of merging iterations under a fixed training budget. With the total training steps fixed (Fig. 6), we compare merging iterations $k\in\{1,4,10\}$. Increasing $k$ consistently improves final performance and reduces degradation from merging multiple experts. For $k=10$ relative to $k=4$ and $k=1$, the gains are +1.12% and +11.84% on GenEval [11], +0.57% and +1.56% on PickScore [15], and +0.03% and +10.72% on OCR [6].
[RaTE] Effect of appended token count. Table 4 reports RaTE performance as we vary the number of appended tokens. In the table, <<RaTE>> denotes the metric-specific token: <<GE>> for GenEval [11], <<PS>> for PickScore [15], and <<OCR>> for OCR [6]. Performance peaks differ by reward: GenEval saturates at 2–3 tokens (0.78), PickScore peaks at 1 (22.052), and OCR peaks at 3 (0.635).
@@ 5 Conclusion
We tackled the core challenge of multi-preference post-training for generative models, where improving one objective (e.g., text–image alignment) often harms others (e.g., aesthetic quality), creating an alignment tax. We introduced two complementary components—Reward-aware Token Embedding (RaTE) and MapReduce LoRA—that convert this trade-off into a favorable, controllable multi-objective tuning strategy. MapReduce LoRA trains per-reward LoRA experts in parallel, merges them with user-controlled weights, folds the merged adapter into the base model, and iterates to steadily advance the Pareto front. RaTE distills each expert into a lightweight learned token embedding that composes at inference, enabling flexible, a posteriori preference control without retraining. Across Text-to-Image (Stable Diffusion 3.5 Medium, FLUX.1-dev), Text-to-Video (HunyuanVideo) and language task (Helpful Assistant, with Llama-2 7B), our approach delivers substantial gains: on T2I, we improve GenEval, PickScore, and OCR by 36.1%, 4.6%, and 55.7% (SD 3.5 M) and 32.7%, 4.3%, and 67.1% (FLUX.1-dev); on T2V, visual and motion quality improve by 48.1% and 90.0%, respectively. On language task, helpful and harmless improve by 43.4% and 136.7%, respectively. Beyond the targeted rewards, untargeted metrics (i.e., VQAScore, MPS, VILA) also increase, indicating robust cross-preference generalization. Overall, MapReduce LoRA offers a simple, scalable recipe for systematically pushing the multi-preference Pareto front and enabling practical post-training customization.
@@ C.1 Training Configuration
      FLUX.1-dev [17]: The only task-dependent change is the KL ratio $\beta$ ($\beta_{\text{GenEval}}{=}0.04$, $\beta_{\text{PickScore}}{=}0$, $\beta_{\text{OCR}}{=}0.04$). The sampling timestep $T$ is 6 and the evaluation timestep $T$ = 28. The GRPO group size $G$ is 24, the resolution is 512, and the LoRA settings are $\alpha$ = 128 and $r$ = 64. LoRA is applied to attention layers and feed-forward network layers (attn.to_q, attn.to_k, attn.to_v, attn.to_out.0, attn.add_q_proj, attn.add_k_proj, attn.add_v_proj, attn.to_add_out, $f$.net.0.proj, $f$.net.2, ff_context.net.0.proj,
      ff_context.net.2). The per-GPU batch size is 3 with gradient accumulation steps of 6 for all three tasks. We train GenEval, PickScore and OCR for {2,700, 1,900, 1,800, 1,600}, {1,250, 600, 1,550, 2,200}, and {1,250, 650, 850, 1,100} steps, respectively; the bracketed numbers denote merge iterations 0-3.
@@ E.1 Text-to-Image qualitative results
Figs 11, 12 and 13 demonstrate the visual comparison of two rewards with different merging ratios. Figs 14 and 15 demonstrate the visual comparison of all methods, including Flow-GRPO [23] tuned on GenEval [11], PickScore [15], OCR [6], Rewarded Soup [33], MORL-D, MORL-DR and MapReduce LoRA. PickScore and OCR guidance cause strong reward overfitting: the former drives the model toward a narrow, high-scoring aesthetic style, while the latter encourages increasingly large and prominent text regardless of overall visual harmony or contextual appropriateness. Our proposed MapReduce LoRA mitigates these issues by iteratively merging per-reward experts, enabling the model to discover a more balanced (and often near-optimal) preference that harmonizes aesthetic fidelity with text readability.
@@ E.2 Text-to-Video qualitative results
Fig. 16 demonstrates the visual comparison of HunyuanVideo [16], Rewarded Soup [33] and our proposed MapReduce LoRA. Also, Figs. 17 and 18 demonstrate the visual performance of merging progress that advances the visual quality, motion quality, and prompt alignment.
@@ Appendix F Limitations and Future Works
  • Scaling to more preferences: In Text-to-Image, we validate on three targeted preferences and three additional untargeted ones; extending to larger numbers of preferences is a promising scaling direction.

### v5

@@ 1 Introduction
To address these limitations, we propose two complementary approaches—MapReduce LoRA and Reward-aware Token Embedding (RaTE)—that jointly advance multi-preference alignment. MapReduce LoRA decomposes multi-objective optimization into i) a Map phase that trains reward-specific LoRA experts in parallel, and ii) a Reduce phase that merges experts with user-controlled interpolation, folds the merged adapter into the base, and iterates to advance the Pareto front. Complementarily, RaTE introduces a lightweight inference-time control by assigning each reward a learned token embedding, enabling composable conditioning by appending multiple tokens to the input prompt. Together, MapReduce LoRA and RaTE enable a posteriori customization without retraining, yielding unified generative models that excel across multiple reward dimensions.
Extensive experiments demonstrate that our proposed methods substantially advance the Pareto front of multi-preference alignment across both Text-to-Image, Text-to-Video generation, and Language Tasks (Fig. 1). On Stable Diffusion 3.5 Medium [3] and FLUX.1-dev [20], our method improves GenEval [14], PickScore [18], and OCR [8] by 36.1%, 4.6%, 55.7% and 32.7%, 4.3%, 67.1%, respectively. Beyond in-domain metrics, untargeted rewards—i.e., VQAScore [24], MPS [47], and VILA [17]—also improve by 1.85%, 6.49%, and 19.96%, validating the robustness and scalability of our method. On HunyuanVideo [19], visual and motion quality improve by 48.1% and 90.0%, achieving state-of-the-art results among post-trained diffusion systems. On Llama-2 7B [42], helpful and harmless improve by 43.4% and 136.7%, respectively. Collectively, these results highlight the effectiveness of combining MapReduce LoRA and RaTE for efficient, controllable, and human-aligned generation. Our contributions are fourfold:
@@ 2.1 Flow-based Generative Models
Flow Matching (FM) [25] trains continuous normalizing flows by regressing the conditional velocity field along a path between the data distribution and a standard normal, yielding a score-free, simulation-free objective with more stable and efficient optimization than diffusion losses. Rectified Flow (RF) [28] specializes FM to a straight-line, approximately constant-speed path, rectifying trajectories and simplifying supervision, which improves stability and sampling efficiency. Recent Text-to-Image [11, 3, 20, 9, 5] and Text-to-Video [19, 34, 43] methods adopt FM/RF to learn the velocity field directly, improving efficiency over diffusion-style denoising and often yielding stronger gradients for conditional generation.
@@ 2.2 Reinforcement Learning from Human Feedback (RLHF)
However, single-reward RLHF optimizes one preference dimension in isolation, while generative quality spans multiple and often conflicting objectives. This creates a natural bridge to Multi-Objective RL (MORL), where the objective is to balance conflicting rewards and approximate a Pareto-optimal solution—precisely the challenge in post-training generative models. MORL methods can be categorized into two classes: a priori [21, 1], which scalarizes objectives into a single signal (e.g., weighted sums), and a posteriori [36, 23], which approximate a Pareto set and enable test-time control. CaPO [21] calibrates and balances multiple T2I preferences, and MOPO [1] frames alignment as constrained, KL-regularized optimization; both are a priori and lack test-time control. Rewarded Soups [36] extends Linear Mode Connectivity [12, 31] to RL, showing linear weight soups can approximate Pareto fronts and reduce reward misspecification, yet multi-reward methods often underperform specialized single-reward experts. We therefore propose two a posteriori solutions, RaTE and MapReduce LoRA, to address reward conflict. Specifically, the iterative merging in MapReduce LoRA shares a technical spirit with periodic averaging in FedAvg [41] and DiLoCo [10]; however, while they focus on communication efficiency across GPU clusters, our work leverages this mechanism to navigate the Pareto front for multi-preference alignment.
@@ 3.1 Preliminaries
where $A\in\mathbb{R}^{r\times k}$, $B\in\mathbb{R}^{d\times r}$, and $r\ll\min(d,k)$. During fine-tuning, $W$ remains fixed while only $A$ and $B$ are trained, optionally scaled by a factor $\alpha/r$; at inference, $\Delta W$ can be merged into $W$. Recent work [37] finds that, in common post-training settings (small/medium-scale SFT, reasoning, or RL), LoRA matches full fine-tuning performance when rank and layer coverage suffice. LoRA has been used for style [48] and skill control [23], but multi-preference control remains underexplored; we therefore study LoRA merging for multi-preference optimization.
@@
Pareto Fronts (PF) [33] represents the collection of non-dominated solutions, model parameters for which no other candidate can improve one objective without degrading another. Formally, $\mathrm{PF}=\left\{\theta\mid\nexists\,\theta^{\prime}\in\Theta,\;\{R_{i}(\theta^{\prime})\}_{i}\succ_{\mathbb{R}^{N}}\{R_{i}(\theta)\}_{i}\right\},$ where $\succ_{\mathbb{R}^{N}}$ denotes the dominance relation in the N-dimensional reward space indexed by $i$, and $N$ denotes the total number of reward objectives.
@@ 4 Experiments
α The detailed settings of MORL-D/DR: Please refer to the Supplement C.
@@ 4.3 Quantitative Results
[Ablation] MapReduce LoRA: Effect of merging iterations under a fixed training budget. With the total training steps fixed (Fig. 6), we compare merging iterations $k\in\{1,4,10\}$. Increasing $k$ consistently improves final performance and reduces degradation from merging multiple experts. For $k=10$ relative to $k=4$ and $k=1$, the gains are +1.12% and +11.84% on GenEval [14], +0.57% and +1.56% on PickScore [18], and +0.03% and +10.72% on OCR [8].
Reward-aware token embedding (RaTE). Tables 1 and 3 report RaTE performance on GenEval [14], PickScore [18], and OCR [8] using SD 3.5 M [3]. At inference, RaTE appends a trained control token to the end of the prompt based on the specified preference. RaTE enables per-reward control and jointly improves multiple rewards. In Table 1, token gains are orthogonal to transformer tuning: RaTE adds 4.20%, 1.03%, and 2.17% on GenEval [14], PickScore [18], and OCR [8], respectively. In Table 3, when using all three tokens, the first appended token accounts for most of the gain. RaTE enables lightweight, composable inference-time control and costs only 0.1579× of MapReduce LoRA to train. Token control is effective on Stable Diffusion [3, 13, 6], which uses explicit cross-attention for text conditioning. In contrast, FLUX.1-dev [20] performs joint text–image sequence modeling, causing a modified token embedding to perturb both text and image tokens across layers, which makes reward-specific token control substantially less stable. We leave this for future exploration.
@@
[Ablation] RaTE: Effect of appended token count. Table 4 reports RaTE performance as we vary the number of appended tokens. In the table, <<RaTE>> denotes the metric-specific token: <<GE>> for GenEval [14], <<PS>> for PickScore [18], and <<OCR>> for OCR [8]. Performance peaks differ by reward: GenEval saturates at 2–3 tokens (0.78), PickScore peaks at 1 (22.052), and OCR peaks at 3 (0.635).
@@ 5 Conclusion
We present MapReduce LoRA and RaTE, two complementary methods to address the ”alignment tax” in multi-preference post-training. By iteratively merging parallel reward experts and distilling them into composable token embeddings, our framework converts objective trade-offs into a controllable tuning strategy. Extensive evaluations across text-to-image (SD 3.5, FLUX.1), video (HunyuanVideo), and language (Llama-2) models demonstrate substantial gains in both targeted and untargeted metrics. Overall, our approach offers a simple, scalable recipe for systematically advancing the multi-preference Pareto front and enabling practical model customization.
@@ C.1 Training Configuration
      FLUX.1-dev [20]: The only task-dependent change is the KL ratio $\beta$ ($\beta_{\text{GenEval}}{=}0.04$, $\beta_{\text{PickScore}}{=}0$, $\beta_{\text{OCR}}{=}0.04$). The sampling timestep $T$ is 6 and the evaluation timestep $T$ = 28. The GRPO group size $G$ is 24, the resolution is 512, and the LoRA settings are $\alpha$ = 128 and $r$ = 64. LoRA is applied to attention layers and feed-forward network layers (attn.to_q, attn.to_k, attn.to_v, attn.to_out.0, attn.add_q_proj, attn.add_k_proj, attn.add_v_proj, attn.to_add_out, ff.net.0.proj, ff.net.2, ff_context.net.0.proj, ff_context.net.2). The per-GPU batch size is 3 with gradient accumulation steps of 6 for all three tasks. We train GenEval, PickScore and OCR for {2,700, 1,900, 1,800, 1,600}, {1,250, 600, 1,550, 2,200}, and {1,250, 650, 850, 1,100} steps, respectively; the bracketed numbers denote merge iterations 0-3.
@@ Appendix D More Quantitative Results
D.3 Scalability of Reward Numbers
We extend MapReduce LoRA to 5 rewards by adding VQAScore [24] and MPS [47] on SD 3.5 M [3]. Fig. 11 shows that MapReduce LoRA improves all 5 rewards (+22% GenEval, +3% PickScore, +44% OCR, +6% VQAScore, +7% MPS), consistently outperforming one-shot Rewarded Soup (the starting point of k = 1). Scaling to more rewards does not introduce diminishing returns: performance on overlapping rewards (GenEval, PickScore, OCR) remains improved compared to 3-reward training (Fig. 6, k=4). However, achieving similar levels requires  1.66× more steps, reflecting increased optimization complexity. Beyond the increased optimization complexity for existing rewards, adding rewards incurs a linear per-reward training cost; however, the parallel Map phase helps mitigate the resulting wall-clock growth given sufficient GPUs.
Figure 11: MapReduce LoRA across 5 rewards with 3 merges.
@@ E.1 Text-to-Image qualitative results
Figs 12, 13 and 14 demonstrate the visual comparison of two rewards with different merging ratios. Figs 15 and 16 demonstrate the visual comparison of all methods, including Flow-GRPO [26] tuned on GenEval [14], PickScore [18], OCR [8], Rewarded Soup [36], MORL-D, MORL-DR and MapReduce LoRA. PickScore and OCR guidance cause strong reward overfitting: the former drives the model toward a narrow, high-scoring aesthetic style, while the latter encourages increasingly large and prominent text regardless of overall visual harmony or contextual appropriateness. Our proposed MapReduce LoRA mitigates these issues by iteratively merging per-reward experts, enabling the model to discover a more balanced (and often near-optimal) preference that harmonizes aesthetic fidelity with text readability.
@@ E.2 Text-to-Video qualitative results
Fig. 17 demonstrates the visual comparison of HunyuanVideo [19], Rewarded Soup [36] and our proposed MapReduce LoRA. Also, Figs. 18 and 19 demonstrate the visual performance of merging progress that advances the visual quality, motion quality, and prompt alignment.
@@ Appendix F Limitations and Future Works
  • Scaling to more preferences: In Text-to-Image, we validate on five targeted preferences; extending to larger numbers of preferences is a promising scaling direction.